### v1

# DiaBlo: Diagonal Blocks Are Sufficient For Finetuning

Abstract
Fine-tuning is a critical step for adapting large language models (LLMs) to domain-specific downstream tasks. To mitigate the substantial computational and memory costs of full-model fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to update only a small subset of model parameters. However, performance gaps between PEFT approaches and full-model fine-tuning still exist. In this work, we present DiaBlo, a simple yet effective PEFT approach that updates only the diagonal blocks of selected model weight matrices. Unlike Low-Rank Adaptation (LoRA) and its variants, DiaBlo eliminates the need for low-rank matrix products, thereby avoiding the reliance on auxiliary initialization schemes or customized optimization strategies to improve convergence. This design leads to stable and robust convergence while maintaining comparable memory efficiency and training speed to LoRA. We conduct extensive experiments across a range of tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—to evaluate the effectiveness and efficiency of DiaBlo. Across these benchmarks, DiaBlo demonstrates strong and consistent performance while maintaining high memory efficiency and fast fine-tuning speed. Codes are available at https://github.com/ziyangjoy/DiaBlo.
1 Introduction
Large language models (LLMs) Achiam2023GPT4TR ; Brown2020LanguageMA ; Touvron2023LLaMAOA have achieved remarkable success across a wide range of natural language processing (NLP) tasks, including reasoning, generation, and alignment. These models, often consisting of billions of parameters, are typically pre-trained on broad, general-purpose corpora. However, to adapt them to specific downstream tasks or domains, fine-tuning is essential. Full fine-tuning, which updates all parameters of the model, has been shown to yield strong performance but is prohibitively expensive in terms of computational cost, memory usage, and storage—especially on resource-constrained devices.
To address these challenges, Parameter-Efficient Fine-Tuning (PEFT) methods Ding2023ParameterefficientFO ; Han2024ParameterEfficientFF have emerged as a promising alternative. These methods aim to retain the performance benefits of full fine-tuning while updating only a small fraction of the model’s parameters. Early approaches, such as Prompt Tuning Lester2021ThePO and Prefix Tuning Li2021PrefixTuningOC , introduce task-specific embeddings or continuous prompts that steer model behavior without modifying its original weights. These methods are lightweight and straightforward to implement, making them suitable for low-resource adaptation scenarios. However, they often struggle with scalability and limited expressiveness, especially on complex tasks. To improve performance, more structured methods like Low-Rank Adaptation (LoRA) Hayou2024LoRAEL ; Hu2021LoRALA have been proposed. LoRA adds trainable low-rank matrices to the pretrained weight, significantly reducing the number of trainable parameters. It has demonstrated strong performance across a variety of applications. Despite its efficiency, LoRA and its variants often encounter issues such as unstable convergence and performance degradation. A growing number of work has proposed various extensions to LoRA, including tailored initialization schemes Meng2024PiSSAPS ; Wang2024MiLoRAHM ; Wang2024LoRAGALA and structure-specific optimization strategies DoRA ; Wang2024LoRAProAL . Nonetheless, such improvements may lead to greater algorithmic complexity, potentially impacting implementation and efficiency.
Figure 1: Comparison between full fine-tuning, LoRA, and proposed DiaBlo.
In this work, we propose DiaBlo, a simple yet effective PEFT framework that fine-tunes only the diagonal blocks of the model’s weight matrices. Unlike LoRA, which introduces low-rank adaptation through the product of two trainable matrices—a structure that often requires careful initialization and customized optimization to ensure stable training—DiaBlo directly updates a structured subset of the model’s original weight parameters. By avoiding the use of matrix products, DiaBlo eliminates the inherent optimization difficulties associated with low-rank decomposition. This leads to greater training stability and more reliable convergence without the need for special tricks or tuning. Despite its simplicity, DiaBlo retains the memory and computational efficiency of LoRA while offering improved representational flexibility. It integrates easily into standard training pipelines and consistently outperforms existing PEFT methods across a broad range of tasks, demonstrating that selectively updating diagonal blocks is a powerful and practical alternative for parameter-efficient fine-tuning. Figure 1 illustrates the difference between full fine-tuning, LoRA, and DiaBlo.
Geometric interpretation of DiaBlo is based on the observation that LoRA’s representational capacity is inherently directionally and globally structured. By introducing a diagonal matrix into the PEFT model, we can apply axis-aligned, per-dimension scaling that LoRA cannot express. Geometrically, this means that while LoRA adjusts the weight matrix by steering it within a constrained low-rank plane, the diagonal component expands this plane into a sheared, anisotropically rescaled region, enabling the model to fine-tune the sensitivity of individual features. This additional flexibility is especially valuable in transfer learning for reasoning tasks, where subtle, independent shifts in feature activation can significantly improve generalization with minimal parameter overhead. Thus, the diagonal augmentation enhances the expressiveness while maintaining high efficiency.
We evaluate DiaBlo across a range of comprehensive tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—to rigorously assess its effectiveness and efficiency. Across all benchmarks, DiaBlo consistently achieves state-of-the-art performance among PEFT approaches, often narrowing or even surpassing the gap to full fine-tuning. These results highlight not only the expressive power of diagonal block updates but also the method’s robustness across diverse domains. DiaBlo demonstrates strong convergence behavior, efficient resource usage, and broad applicability, making it a compelling alternative to existing parameter-efficient methods.
Our contributions are summarized as follows:
  • A novel PEFT framework: We propose DiaBlo, a simple yet effective method that fine-tunes only the diagonal blocks of model weight matrices. This structured approach enables efficient adaptation while achieving strong performance across a variety of downstream tasks.
  • Stable optimization behavior: DiaBlo directly updates diagonal blocks using standard training pipelines, without relying on matrix product structures or specialized initialization, resulting in more stable and reliable fine-tuning.
  • Strong empirical performance: Through extensive experiments on tasks including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment, we demonstrate that DiaBlo consistently achieves strong downstream performance, often surpassing LoRA and its variants in accuracy.
  • Memory and computational efficiency: DiaBlo retains the low memory footprint and fast training speed of LoRA and other PEFT methods, ensuring practical usage for large-scale fine-tuning.
2 Background and Related Works
Early PEFT techniques, such as Prompt Tuning Han2021PTRPT ; Lester2021ThePO and Prefix Tuning Li2021PrefixTuningOC , introduce learnable embeddings or token sequences that steer the model’s behavior without modifying the core parameters. These methods are lightweight and easy to integrate, offering low computational overhead. While effective in low-resource or few-shot scenarios, prompt-based methods often struggle to match the performance of full fine-tuning on more complex or general-purpose tasks due to their limited expressiveness and adaptability. Adapter-based approaches Houlsby2019ParameterEfficientTL ; Rckl2020AdapterDropOT ; Wang2020KAdapterIK insert small, trainable bottleneck layers into the transformer architecture. These modules enable task-specific learning while keeping most of the model’s parameters frozen. Variants of adapter tuning further improve performance through mechanisms such as attention fusion or parameter sharing. Although adapters may reach a good performance, they often require changes to model architecture and introduce additional inference-time latency. In addition, there exist matrix or tensor decomposition-based approaches Lingam2024SVFTPF ; Liu2025CoLACP ; Wang2024SVDLLMTS ; Yang2024LoRETTALE ; Yang2024CoMERACA ; Yang2023QuantizationAwareAT to reduce training or finetuning costs.
Low-Rank Adaptation and Variants
Low-Rank Adaptation (LoRA) Hu2021LoRALA has emerged as one of the most popular PEFT methods. LoRA injects low-rank matrices into the update paths of existing model weights, significantly reducing the number of trainable parameters. Its effectiveness and simplicity have led to wide adoption, but its core reliance on the product of two low-rank matrices introduces optimization challenges. This structure can restrict representational flexibility and make training sensitive to initialization and optimization algorithms.
Several recent variants have been proposed to enhance the performance of LoRA. DoRA DoRA introduces a novel optimization strategy that decouples the magnitude and direction of low-rank updates, leading to more stable training dynamics and better performance. Pissa Meng2024PiSSAPS and MiLoRA Wang2024MiLoRAHM focus on different initialization techniques for low-rank adapters. Pissa initializes the low-rank matrices using the largest singular values and singular vectors of weights, thereby preserving the most significant components of the weight matrix for faster convergence. On the other hand, MiLoRA takes a different approach by utilizing the smallest singular values during initialization. This strategy captures the finer details of the weight matrix. Additionally, LoRA-GA Wang2024LoRAGALA proposes to initialize low-rank matrices to align with gradients of full fine-tuning at the first step. While these LoRA-based methods have achieved notable improvements, they often introduce additional complexity in design and implementation. Furthermore, the dependence on matrix-product-based updates continues to present challenges for stable and efficient optimization. There exist many other variants of LoRA. We refer readers to Kalajdzievski2023ARS ; Valipour2022DyLoRAPT ; Wang2024RoseLoRARA ; Zhang2023AdaLoRAAB ; Zhong2024NEATNP . QLoRA dettmers2023qlora reduces the memory footprint of fine-tuning by applying low-rank adaptation on quantized models. Building on this idea, subsequent works GPTQLoRA ; deng2025cloq ; li2023loftq ; liao2024apiq design customized quantization algorithms and develop effective initialization strategies for the low-rank adaptation components, aiming to improve fine-tuning performance.
3 Methodology
3.1 DiaBlo: Finetuning Diagonal Blocks
Motivation
Despite the success of LoRA-based fine-tuning methods, low-rank factorizations remain difficult to optimize in practice. In computational mathematics, problems such as low-rank approximation and low-rank matrix completion are well known to be non-convex and challenging Ma2019BeyondPB ; Tong2020AcceleratingIL ; Tu2015LowrankSO . These difficulties are reflected in LoRA-based approaches, where the product of two trainable low-rank matrices can lead to unstable gradient flow and affect convergence. As a result, much recent work has focused on designing better initialization schemes and custom optimization strategies to stabilize training. The development of DiaBlo is motivated by these limitations. To this end, DiaBlo directly fine-tunes only the diagonal blocks of model weight matrices, avoiding matrix-product parameterizations entirely. This design simplifies the optimization landscape, improves training stability, and enables effective adaptation without introducing additional complexity.
Algorithm
The state-of-the-art LLMs are constructed based on linear layers. Hence, it is sufficient to present our proposed approach on a simple linear layer. Consider a linear layer with input features $m_{1}$ and output features $m_{2}$,
$\mathbf{Y}=\mathbf{X}\mathbf{W},$
where $\mathbf{Y}\in\mathbb{R}^{b\times m_{2}},\mathbf{X}\in\mathbb{R}^{b\times m_{1}},\mathbf{W}\in\mathbb{R}^{m_{1}\times m_{2}}$, and $b$ is the batch size. For a given integer $N$ that is a factor of both $m_{1},m_{2}$, we rewrite the weight $\mathbf{W}$ in the form of block matrices
$\mathbf{W}=\begin{pmatrix}\mathbf{W}_{11}&\mathbf{W}_{12}&\cdots&\mathbf{W}_{1N}\\ \mathbf{W}_{21}&\mathbf{W}_{22}&\cdots&\mathbf{W}_{2N}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{W}_{N1}&\mathbf{W}_{N2}&\cdots&\mathbf{W}_{NN}\\ \end{pmatrix},$
where $\mathbf{W}_{ij}\in\mathbb{R}^{d_{1}\times d_{2}},d_{1}=\frac{m_{1}}{N},d_{2}=\frac{m_{2}}{N}$. During finetuning, only the diagonal blocks of $\mathbf{W}$, i.e., $\mathbf{W}_{11},\ldots,\mathbf{W}_{NN}$, are trainable while all other blocks $\{\mathbf{W}_{ij}\}_{i\neq j}$ are frozen. DiaBlo also works when $N$ is not a common factor of $m_{1},m_{2}$. In that case, we may expand the pad the weight $\mathbf{W}$ into a proper size and then select the corresponding diagonal blocks.
Efficient Implementation
For convenience of implementation, we use a similar adaptation method like LoRA and its variants. We introduce a block diagonal adaptation matrix $\mathbf{D}$ to the linear layer:
$\mathbf{Y}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{D}.$ (1)
The adaptation matrix $\mathbf{D}$ is block diagonal
$\mathbf{D}=\begin{pmatrix}\mathbf{D}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{D}_{2}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{D}_{N}\\ \end{pmatrix},$
where $\mathbf{D}_{i}\in\mathbb{R}^{d_{1}\times d_{2}}$. During fine-tuning, only $\mathbf{D}_{i}$’s are trainable. In real implementations, the adaptation matrix $\mathbf{D}$ can be saved as a tensor $\mathbf{\mathcal{D}}\in\mathbb{R}^{N\times d_{1}\times d_{2}}$, where $\mathbf{\mathcal{D}}_{i,:,:}=\mathbf{D}_{i}$. During forward- and back- propagation, it is not necessary to reconstruct the matrix $\mathbf{D}$ as well. The matrix product $\mathbf{X}\mathbf{D}$ can be written as $\begin{pmatrix}\mathbf{X}_{1}\mathbf{D}_{1}&\cdots&\mathbf{X}_{N}\mathbf{D}_{N}\end{pmatrix},$ where $\mathbf{X}_{i}\in\mathbb{R}^{b\times d_{1}}$. Therefore, $\mathbf{X}\mathbf{D}$ is mathematically equivalent to batched matrix multiplications, which can be computed efficiently on GPUs. In PyTorch, the operation can be easily implemented using torch.einsum. After reshaping $\mathbf{X}$ into tensor $\mathbf{X}$ of size $b\times N\times d_{1}$, $\mathbf{X}\mathbf{D}$ can be computed as
[…]
Let $\mathbf{g}_{\mathbf{Y}}$ be the gradient of loss to the output $\mathbf{Y}$, then the gradient to $\mathbf{D}_{i}$ is
$\mathbf{g}_{\mathbf{D}_{i}}=\mathbf{X}_{i}^{\top}\mathbf{g}_{\mathbf{Y}_{i}},$ (2)
where $\mathbf{g}_{\mathbf{Y}_{i}}=(\mathbf{g}_{\mathbf{Y}})_{:,(i-1)d_{2}+1,id_{2}}$. Hence, the backpropagation of DiaBlo can be efficiently implemented using batched matrix multiplications as well.
Initialization of DiaBlo
LoRA and its variants typically represent adaptation as a low-rank product $\mathbf{A}\mathbf{B}$. Standard initialization schemes set $\mathbf{A}$ using Kaiming uniform and $\mathbf{B}$ as zero, ensuring $\mathbf{A}\mathbf{B}=0$ while the gradients to $\mathbf{A},\mathbf{B}$ do not vanish. However, such initialization often leads to suboptimal performance, as observed in recent studies Meng2024PiSSAPS ; Wang2024MiLoRAHM ; Wang2024LoRAGALA . Consequently, variants like Pissa Meng2024PiSSAPS , MiLoRA Wang2024MiLoRAHM , and LoRA-GA Wang2024LoRAGALA propose alternative initialization techniques to improve optimization behavior—e.g., Pissa uses large singular values, while MiLoRA leverages small singular values for initialization. In contrast, DiaBlo does not involve matrix products and can be initialized more straightforwardly. We simply initialize $\mathbf{\mathcal{D}}$ as an all-zero tensor. This avoids issues related to vanishing gradients or entangled parameter updates, enabling stable and efficient training without the need for specialized initialization and optimization techniques.
3.2 Comparison with LoRA
Gradient Analysis
The gradient with respect to the full weight matrix $\mathbf{W}$ is given by:
$\mathbf{g}_{\mathbf{W}}=\mathbf{X}^{\top}\mathbf{g}_{Y}=\begin{pmatrix}\mathbf{X}_{1}^{\top}\\ \vdots\\ \mathbf{X}_{N}^{\top}\end{pmatrix}\begin{pmatrix}\mathbf{g}_{Y_{1}}&\cdots&\mathbf{g}_{Y_{N}}\end{pmatrix}=\begin{pmatrix}\mathbf{X}_{1}^{\top}\mathbf{g}_{Y_{1}}&\cdots&\mathbf{X}_{1}^{\top}\mathbf{g}_{Y_{N}}\\ \vdots&\ddots&\vdots\\ \mathbf{X}_{N}^{\top}\mathbf{g}_{Y_{1}}&\cdots&\mathbf{X}_{N}^{\top}\mathbf{g}_{Y_{N}}\\ \end{pmatrix}.$
This implies that the gradient of the DiaBlo adaptation block $\mathbf{D}_{i}$ is exactly the gradient of the corresponding diagonal block $\mathbf{W}_{ii}$, i.e., $\mathbf{g}_{\mathbf{D}_{i}}=\mathbf{g}_{\mathbf{W}_{ii}}$. Therefore, optimizing DiaBlo’s diagonal block adaptation is mathematically equivalent to fine-tuning the diagonal blocks of $\mathbf{W}$, which closely mirrors the behavior of full-model fine-tuning in those subspaces.
In contrast, for LoRA, where the adaptation takes the form $\mathbf{Y}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{A}\mathbf{B}$, the gradients with respect to the low-rank matrices are computed as:
$\mathbf{g}_{\mathbf{A}}=\mathbf{g}_{\mathbf{W}}\mathbf{B}^{\top},\quad\mathbf{g}_{\mathbf{B}}=\mathbf{A}^{\top}\mathbf{g}_{\mathbf{W}}.$
These gradients depend on the values of both $\mathbf{A}$ and $\mathbf{B}$, and can suffer from vanishing or unstable updates. As a result, LoRA’s performance is sensitive to initialization strategies and often requires carefully designed optimization techniques. In contrast, DiaBlo directly applies gradients to the full-rank diagonal blocks, making it inherently more stable and effective for fine-tuning.
Low-Rank vs. Full-Rank
Although the block diagonal adaptation $\mathbf{D}$ in DiaBlo is initialized to zero, it is not constrained to be low-rank during training. In contrast to LoRA, which limits updates to a low-rank subspace, DiaBlo operates in a restricted full-rank subspace formed by the diagonal blocks of the weight matrix. While it is theoretically difficult to determine which subspace—low-rank or diagonal-block—is universally more expressive or optimal for fine-tuning, our empirical results suggest that the diagonal-block subspace used by DiaBlo is sufficiently expressive for a wide range of downstream tasks. In practice, DiaBlo consistently delivers strong performance, often surpassing LoRA in accuracy while maintaining similar efficiency.
4 Experiments
In this section, we evaluate the performance of DiaBlo on a variety of benchmark datasets. Specifically, we assess its effectiveness on commonsense reasoning tasks, arithmetic reasoning tasks, code generation tasks, and safety alignment tasks. Our experiments are conducted using LLaMA2-7B touvron2023llama2 , LLaMA3-8B Dubey2024TheL3 , and Mistral-7B-v0.1 Jiang2023Mistral7 models. All fine-tuning and evaluation are performed on a single NVIDIA A100 GPU with 80GB of memory. The implementation details are provided in Appendix A.
4.1 Commonsense Reasoning
Models and Datasets
Commonsense reasoning 170k includes eight benchmark datasets: BoolQ clark2019boolq , PIQA bisk2020piqa , SocialIQA sap2019socialiqa , HellaSwag zellers2019hellaswag , WinoGrande sakaguchi2021winogrande , ARC-Easy, ARC-Challenge clark2018think , and OpenBookQA mihaylov2018can . Each task is framed as a multiple-choice problem. We fine-tune the LLaMA2-7B and LLaMA3-8B models on the combined training data from all eight tasks, and report accuracy on the test set of each task individually.
Implementation Details
Our fine-tuning and evaluation follow the standard settings in hu-etal-2023-llm ; DoRA . Specifically, we fine-tune the Q, K, V, Up, and Down modules using a maximum sequence length of 256. The model is loaded in float16 precision, the DiaBlo adaptation is in float32 precision, and FP16 mixed precision training is enabled.
Results
The results on Commonsense reasoning tasks are shown in Table 1. We compare DiaBlo with existing methods, including DoRA DoRA , Pissa Meng2024PiSSAPS , MiLoRA Wang2024MiLoRAHM , and SMT He2024SparseMI . DiaBlo consistently outperforms existing PEFT baselines on commonsense reasoning tasks using both Llama2-7B and Llama3-8B models in FP16 precision. For Llama2-7B, DiaBlo achieves the highest average score of 83.5% while using only 0.52% trainable parameters, surpassing all other methods with similar number of parameters by at least 1.7%. The SMT can reach an average of 83.4% but requires 4.91% trainable parameters. For Llama3-8B, DiaBlo with $N=64$ and 1.04% trainable parameters achieves the best overall average score of 87.3%, surpassing LoRA and DoRA by 6.5% and 2.1%, respectively. The best SMT variant achieves 87.2%, which uses significantly more parameters. DiaBlo also demonstrates strong and consistent results across individual tasks. The findings demonstrate DiaBlo is an efficient and effective fine-tuning strategy that offers competitive accuracy while using few trainable parameters.
Comparison with SMT
SMT He2024SparseMI proposes to select most important sub-matrices in pretrained weights based on magnitudes of gradients and then only fine-tune these sub-matrices. In contrast, DiaBlo directly selects diagonal blocks from the pretrained weights, offering a simpler and more implementation-friendly approach. The experiments further demonstrate that fine-tuning diagonal blocks provide even better results. As shown in Table 1, DiaBlo achieves an average score of 83.5% on Llama2-7B, significantly outperforming SMT’s 81.8% while using a less number of trainable parameters. Even when SMT increases its trainable parameter to 4.91% to achieve its best result, DiaBlo, with only 0.52% trainable parameters, still attains better performance. A similar trend holds for Llama3-8B, where DiaBlo outperforms SMT even when SMT uses considerably more trainable parameters. These results demonstrate that fine-tuning only the diagonal blocks of pretrained weights is not only highly efficient but also sufficient to achieve strong performance.
Precision
The experiments in existing literature, e.g., DoRA DoRA , Pissa Meng2024PiSSAPS , and MiLoRA Wang2024MiLoRAHM , all follow the fine-tuning setting in hu-etal-2023-llm , which uses FP16 for model weights, FP32 for low-rank adaptation, and FP16 mixed-precision training. Our results reported in Table 1 follow the same fine-tuning setting. However, we notice that LoRA finetuning in precision BF16 may achieve better results. For comparison, we also report results trained in BF16 precision following the setting in LoRI Zhang2025LoRIRC . The results are reported in Table 2. On LLaMA3-8B, DiaBlo achieves the highest average score of 88.3% with $N=64$ blocks and 1.51% of trainable parameters, surpassing LoRA, DoRA, and LoRI by at least 1.0%. Even when reducing the number of trainable parameters to 0.76% with $N=128$ blocks, DiaBlo maintains strong performance, achieving 88.0%, again outperforming the baselines. On Mistral-7B, DiaBlo achieves an average accuracy of 88.0% with $N=64$ blocks, exceeding LoRA, DoRA, and LoRI. At $N=128$ blocks with less trainable parameters, it continues to perform competitively, reaching 87.9%. These results confirm that DiaBlo is highly effective under BF16 precision, offering both superior accuracy and parameter efficiency compared to other state-of-the-art PEFT methods.
Table 1: Commonsense reasoning results of fine-tuning Llama2-7B and Llama3-8B in FP16. Results marked with , , are taken from Dora DoRA , MiLoRA Wang2024MiLoRAHM , and SMT He2024SparseMI , respectively.
Llama2-7B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 0.83% 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
DoRA $r=$32 0.83% 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4 79.7
DoRA $r=$16 0.42% 72.0 83.1 79.9 89.1 83.0 84.5 71.0 81.2 80.5
Pissa $r=$32 0.83% 67.6 78.1 78.4 76.6 78.0 75.8 60.2 75.6 73.8
MiLoRA $r=$32 0.83% 67.6 83.8 80.1 88.2 82.0 82.8 68.8 80.6 79.2
SMT N/A 0.84% 72.0 83.8 80.8 93.3 82.8 86.7 74.0 81.0 81.8
SMT(Best) N/A 4.91% 72.6 85.2 82.0 94.4 85.7 87.8 74.5 85.0 83.4
DiaBlo $N=$64 1.04% 73.9 84.8 81.7 90.0 85.0 87.9 76.8 86.8 83.4
DiaBlo $N=$128 0.52% 74.8 85.5 80.9 89.9 85.7 88.5 76.3 86 83.5
Llama3-8B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 0.78% 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
DoRA $r=$32 0.78% 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
DoRA $r=$16 0.39% 74.5 88.8 80.3 95.5 84.7 90.1 79.1 87.2 85.0
Pissa $r=$32 0.78% 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRA $r=$32 0.78% 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
SMT N/A 0.71% 75.7 88.4 81.4 96.2 88.2 92.7 83.2 88.6 86.8
SMT(best) N/A 3.01% 75.1 89.9 82.4 96.3 88.8 92.6 82.8 89.6 87.2
DiaBlo $N=$64 1.04% 75.8 90.8 80.7 95.6 89.9 93.4 83 89.2 87.3
DiaBlo $N=$128 0.52% 76.1 90.3 81.5 95.7 89.7 93.4 83.4 87.2 87.2
Table 2: Commonsense reasoning results of fine-tuning Llama3-8B and Mistral-7B in BF16 following LoRI. Results of baseline methods are taken from LoRI Zhang2025LoRIRC .
Mistral-7B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 1.25% 75.2 90.1 82.9 95.1 88.1 92.0 82.9 88.7 86.9
DoRA $r=$32 1.25% 75.8 90.4 82.9 96.3 87.9 92.6 83.3 90.6 87.5
LoRI $r=32$ 0.63% 75.9 90.6 83.0 95.9 87.4 91.9 83.6 88.4 87.1
DiaBlo $N=$64 1.68% 76.6 90.9 83.7 96.9 88.6 93.8 83.2 90.5 88.0
DiaBlo $N=$128 0.84% 76.7 91.3 83.9 96.9 88.2 93.1 83.4 89.7 87.9
Llama3-8B
PEFT $\mathbf{r/N}$Params BoolQ PIQA SIQA HellaS WinoG ARC-e ARC-c OBQA AVG
LoRA $r=$32 1.12% 76.3 89.8 82.7 95.8 88.7 91.7 83.4 88.4 87.1
DoRA $r=$32 1.12% 75.9 89.8 82.7 95.3 88.2 93.2 83.5 87.9 87.1
LoRI $r=32$ 0.56% 76.4 89.0 82.7 95.9 87.9 93.6 84.2 88.5 87.3
DiaBlo $N=$64 1.51% 76.6 90.2 83.1 96.9 88.8 93.9 85.1 91.5 88.3
DiaBlo $N=$128 0.76% 76.2 90.5 83.7 97.0 89.0 94.0 84.7 88.9 88.0
4.2 Arithmetic Reasoning
Models and Datasets
We evaluate DiaBlo on arithmetic reasoning using the MetaMathQA Yu2023MetaMathBY dataset, which consists of 395K samples augmented from the training sets of GSM8K cobbe2021training and MATH Hendrycks2021MeasuringMP . We fine-tune the LLaMA2-7B model on this dataset and evaluate performance on the test sets of GSM8K and MATH. The training follows the setting in MiLoRA and we report results from the final checkpoint and use the Exact Match (EM) ratio to measure accuracy.
Results
On arithmetic reasoning tasks, DiaBlo demonstrates strong performance, achieving results that match or exceed full fine-tuning while maintaining parameter efficiency. With $N=32$ blocks and only 2.09% trainable parameters, DiaBlo attains an average accuracy of 43.4%, slightly outperforming full fine-tuning with 43.2% and significantly surpassing LoRA, PiSSA, and MiLoRA by 4.7%, 6.4%, and 2.7%, respectively. Even with $N=64$ blocks and 1.04% trainable parameters, DiaBlo with fewer trainable parameters maintains high performance with an average accuracy of 42.1%, demonstrating that it remains effective under restrictive parameter budgets. Notably, DiaBlo achieves the highest score of 20.4% on the MATH dataset among all methods, including full fine-tuning. These results highlight DiaBlo’s ability to generalize to complex reasoning tasks while offering superior parameter efficiency.
Table 3: Performance of fine-tuning Llama2-7B on arithmetic reasoning tasks. Results of baseline methods are taken from MiLoRA Wang2024MiLoRAHM .
Method $\mathbf{r}/\mathbf{N}$Params GSM8K MATH AVG
Full FT N/A 100% 66.5 19.8 43.2
LoRA $r=64$ 1.67% 60.6 16.9 38.7
PiSSA $r=64$ 1.67% 58.2 15.8 37.0
MiLoRA $r=64$ 1.67% 63.5 17.8 40.7
DiaBlo $N=32$ 2.09% 66.3 20.4 43.4
DiaBlo $N=64$ 1.04% 66.5 17.6 42.1
4.3 Code Generation and Safety Alignment
Models and Datasets
We further evaluate the performance of DiaBlo on code generation and safety alignment tasks, following the training setting in LoRI Zhang2025LoRIRC . For code generation, we fine-tune on the CodeAlpaca dataset codealpaca and evaluate using the HumanEval benchmark Chen2021EvaluatingLL with standard metrics: Pass@1, Pass@5, and Pass@10. For safety alignment, we fine-tune on the SaferPaca dataset Bianchi2023SafetyTunedLL , an extension of Alpaca-Cleaned alpaca , and assess performance based on the refusal rate to harmful prompts from the HEx-PHI dataset Qi2023FinetuningAL . We conduct experiments on both the LLaMA3-8B and Mistral-7B models.
Results
The experimental results on HumanEval demonstrate that DiaBlo achieves strong performance in code generation on Llama3-8B. With 1.51% trainable parameters, DiaBlo reaches a Pass@1 score of 41.8% and a Pass@10 score of 62.9%, outperforming LoRA and DoRA across all metrics and closely matching the best-performing method, LoRI. With a reduced parameter budget of 0.76%, DiaBlo still maintains high performance, achieving a Pass@10 score of 61.9%. Notably, if trained for one additional epoch, DiaBlo further improves to a Pass@1 score of 42.65%, Pass@5 of 57.11%, and Pass@10 of 64.07%, which are comparable to the best results achieved by LoRI. On Mistral-7B, DiaBlo with $N=128$ achieves the best Pass@5 and Pass@10 scores and the second-best Pass@1 score, while DiaBlo with $N=64$ achieves the best Pass@1 score and the second-best Pass@5 and Pass@10 scores, outperforming LoRA, DoRA, and LoRI. In terms of HEx-PHI, DiaBlo outperforms all other methods with scores of 97.6% on Llama3-8B and 98.8% on Mistral-7B. These findings highlight DiaBlo’s effectiveness in both code generation and safety alignment, while maintaining a high level of parameter efficiency.
Table 4: Performance on code generation and safety alignment tasks. Results of baseline methods are taken from LoRI Zhang2025LoRIRC .
Model Method $\mathbf{r}/\mathbf{N}$Params HumanEval HEx-PHI
Pass@1 Pass@5 Pass@10
Llama3-8B LoRA $r=32$ 1.12% 34.7 46.4 50.8 91.6
DoRA $r=32$ 1.12% 33.1 44.0 48.6 93.6
LoRI $r=32$ 0.56% 43.2 57.6 63.2 92.8
DiaBlo $N=64$ 1.51% 41.8 56.4 62.9 97.6
DiaBlo $N=128$ 0.76% 39.4 55.0 61.9 96.3
Mistral-7B LoRA $r=32$ 1.25% 33.8 42.4 45.3 91.9
DoRA $r=32$ 1.25% 33.7 42.6 46.8 95.3
LoRI $r=32$ 0.63% 33.8 42.0 45.1 94.7
DiaBlo $N=64$ 1.68% 34.4 44.8 48.7 98.8
DiaBlo $N=128$ 0.84% 34.0 45.7 49.1 98.8
4.4 Fine-tuning with Quantized Models
To further reduce the memory footprint of fine-tuning, QLoRA guo2024lqlora and its variants store the frozen pretrained weights in low-precision while training small low-rank adaptation modules in higher precision. In this section, we explore the compatibility of the proposed DiaBlo with quantized models and evaluate its performance when fine-tuning with quantized models.
Models and Datasets
We use DiaBlo to fine-tune Llama2-7B model in 4-bit and 2-bit, which are quantized by MagR zhang2024magr . Following the setting in ApiQ liao2024apiq , we fine-tune the model on Math10K hu-etal-2023-llm and then evaluate on test sets of AQuA ling2017program , GSM8K cobbe2021training , MAWPS koncel2016mawps , and SVAMP patel2021nlp .
Results
We evaluate DiaBlo on 4-bit and 2-bit quantized models for arithmetic reasoning tasks and compare its performance with existing fine-tuning methods for quantized models, QLoRA dettmers2023qlora , GPTQ-LoRA GPTQLoRA , LoftQ li2023loftq , and ApiQ liao2024apiq . In the 4-bit setting, DiaBlo achieves the highest average accuracy of 54.8%, outperforming all other methods while also using significantly fewer trainable parameters. In the more challenging 2-bit setting, DiaBlo again outperforms existing baselines with an average accuracy of 48.7%. These results demonstrate that DiaBlo remains robust and effective under low-precision constraints, offering a highly parameter-efficient and accurate approach for fine-tuning quantized models. Notably, DiaBlo directly fine-tunes quantized models without relying on any customized quantization or initialization strategy. In contrast, state-of-the-art methods such as LoftQ li2023loftq and ApiQ liao2024apiq introduce specialized mechanisms to calibrate quantized weights and carefully initialize LoRA components to improve stability and performance. The simplicity of DiaBlo highlights its practical advantages while achieving superior or comparable results.
Table 5: Performance of fine-tuning quantized Llama2-7B model on arithmetic reasoning tasks. Results of baseline methods are taken from ApiQ liao2024apiq .
Method $\mathbf{r/N}$Params GSM8K SVAMP mawps AQuA AVG
LoRA $r=64$ 112M 43.6 59.4 85.0 27.0 53.7
4-bit QLoRA $r=64$ 112M 42.7 58.7 87.3 26.4 53.7
GPTQ-LoRA $r=64$ 112M 43.0 58.4 86.1 24.3 52.9
LoftQ $r=64$ 112M 41.7 56.0 86.3 25.3 52.3
ApiQ-bw $r=64$ 112M 43.2 59.0 85.7 26.0 53.5
DiaBlo $N=64$ 70M 44.1 59.0 89.5 26.4 54.8
2-bit QLoRA $r=64$ 112M 0.9 1.5 0.8 5.1 2.1
GPTQ-LoRA $r=64$ 112M 21.7 39.0 76.6 22.1 39.9
LoftQ $r=64$ 112M 29.5 45.8 83.6 23.2 45.6
ApiQ-bw $r=64$ 112M 31.2 49.0 83.9 23.9 47.3
DiaBlo $N=64$ 70M 32.1 51.5 87.0 24.0 48.7
4.5 Memory and Computation Efficiency of DiaBlo
The simple structure of DiaBlo allows it to make use of efficient batched operations, helping reduce training time and memory usage. In this subsection, we evaluate the memory consumption and training speed of DiaBlo and compare them with other PEFT training methods.
Consider a linear layer $\mathbf{Y}=\mathbf{X}\mathbf{W}$, where $\mathbf{W}\in\mathbb{R}^{m\times m},\mathbf{X}\in\mathbb{R}^{b\times m}$. Let $N$ be the number of blocks in DiaBlo. Each diagonal block is of size $d\times d$, where $d=\frac{m}{N}$, and the total number of trainable parameters is $Nd^{2}$. The forward pass using the block matrix structure requires operations $bNd^{2}$. In comparison, LoRA with rank $r$ has $2mr$ trainable parameters and requires $2bmr$ operations during the forward pass. When DiaBlo and LoRA are configured to have the same number of trainable parameters, that is, $Nd^{2}=2mr$, it is straightforward to show that the computational cost of the forward pass is also the same. A similar argument applies to the backward pass. Therefore, DiaBlo has the same theoretical computational complexity and memory footprint as LoRA when they share the same number of trainable parameters. This indicates that DiaBlo is as efficient as vanilla LoRA in terms of resource usage and computation complexity.
Table 6: Comparison of fine-tuning runtime.
Method $\mathbf{r/N}$Params Time (min)
LoRA $r=32$ 1.12% 170
DoRA $r=32$ 1.12% 480
DiaBlo $N=64$ 1.51% 170
DiaBlo $N=128$ 0.76% 170
To further validate this, we benchmark the fine-tuning speed of DiaBlo on the commonsense reasoning 170k dataset and compare it with LoRA and DoRA. The experiments are conducted on a single A100 GPU with 80GB memory, using a batch size of 8 and a gradient accumulation step of 4. The times for one-epoch fine-tuning are presented in Table 6. DiaBlo achieves training speeds exactly the same as LoRA, while DoRA takes significantly longer. This demonstrates the practical efficiency of the proposed DiaBlo method.
5 Conclusion
In this work, we introduced DiaBlo, a simple yet effective parameter-efficient fine-tuning (PEFT) framework that directly updates diagonal blocks of model weight matrices. Unlike low-rank approaches such as LoRA and its many variants, DiaBlo avoids matrix-product-based adaptations, eliminating the need for special initialization or customized optimization techniques. Through extensive experiments across a diverse set of tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—we demonstrated that DiaBlo consistently achieves strong or state-of-the-art performance while maintaining a comparable or smaller number of trainable parameters. DiaBlo not only matches the theoretical memory and computation efficiency of LoRA, but also shows competitive training speed in practice. Moreover, DiaBlo generalizes well to low-precision settings: in both 4-bit and 2-bit quantized models, it outperforms existing PEFT methods for quantized models without relying on carefully designed quantization or initialization strategies. This robustness and simplicity make DiaBlo an appealing choice for scalable and reliable fine-tuning of large language models. In conclusion, DiaBlo offers a practical and effective approach for efficient model fine-tuning. As large language models and other AI models continue to scale, the demand for robust and straightforward fine-tuning techniques becomes increasingly important. DiaBlo addresses this demand by achieving a strong balance between performance, efficiency, and implementation simplicity. In the future, we will explore the application of DiaBlo to more modalities other than large language models and investigate combinations of DiaBlo with other PEFT approaches, as these methods can be complementary to each other.
6 Limitations
This work focuses on evaluating the fine-tuning performance of DiaBlo on language models. Due to time and computational constraints, we have not explored its applicability to models in other modalities, such as vision or vision-language models. Investigating its potential in those domains remains an important direction for future research. While our empirical results demonstrate that selecting diagonal blocks for fine-tuning yields strong performance on language tasks, the theoretical foundations underlying this effectiveness are not yet well understood. Further exploration into the theoretical properties of diagonal adaptation could provide valuable insights and guide future developments.
Appendix A Implementation Details
A.1 Commonsense Reasoning
The results in Table 1 on commonsense reasoning in FP16 follow the standard setting in [23, 33]. The results in Table 2 on commonsense reasoning in BF16 follow the setting in [61]. The hyper-parameteres for FP16 are provided in 7 and the hyper-parameteres for FP16 are provided in 8
Table 7: Hyper-parameters for the finetuning of Llama2-7B and Llama3-8B on commonsense reasoning in FP16.
Hyper-parameter Llama2-7B Llama3-8B
number of blocks 64 128 64 128
Optimizer AdamW
Learning rate 2e-4 3e-4 5e-5 1e-4
Weight decay 0
LR scheduler linear
Warmup steps 100
Epochs 3
Batch size 16
Max sequence length 256
Modules Q K V U D
Table 8: Hyper-parameters for the finetuning of Mistral-7B and Llama3-8B on commonsense reasoning in BF16.
Hyper-parameter Mistral-7B Llama3-8B
number of blocks 64 128 64 128
Optimizer AdamW
Learning rate 5e-5 5e-5 1e-4 2e-4
Weight decay 0
LR scheduler linear
Warmup steps 100
Epochs 1
Batch size 32
Max sequence length 512
Modules Q K V O G U D
A.2 Arithmetic Reasoning
We follow the training setting of MiLoRA [49] for the arithmetic reasoning task. The hyper-parameter setting for fine-tuning Llama2-7B on MetaMathQA dataset is shown in Table 9.
Table 9: Hyper-parameters for the finetuning of Llama2-7B on MetaMathQA.
Hyper-parameter Llama2-7B
number of blocks 32 64
Optimizer AdamW
Learning rate 2e-4
Weight decay 0
LR scheduler linear
Warmup steps 100
Epochs 3
Batch size 16
Max sequence length 2048
Modules Q K V U D
A.3 Code Generation and Safety Alignment
We follow the training setting of LoRI [61] for the code generation and safety alignment tasks. The hyper-parameter setting for code generation task is shown in Table 10 and the setting for the safety alignment is show in Table 11.
Table 10: Hyper-parameters for the finetuning of Mistral-7B and Llama3-8B on CodeAlpaca dataset.
Hyper-parameter Mistral-7B Llama3-8B
number of blocks 64 128 64 128
Optimizer AdamW
Learning rate 1e-4 1e-4 2e-5 5e-5
Weight decay 0
LR scheduler linear constant linear
Warmup steps 100
Epochs 2
Batch size 32
Max sequence length 512
Modules Q K V O G U D
Table 11: Hyper-parameters for the finetuning of Mistral-7B and Llama3-8B on Saferpaca dataset.
Hyper-parameter Mistral-7B Llama3-8B
number of blocks 64 128 64 128
Optimizer AdamW
Learning rate 1e-4
Weight decay 0
LR scheduler linear
Warmup steps 100
Epochs 1
Batch size 32
Max sequence length 512
Modules Q K V O G U D
A.4 Quantized Models
We finetune with quantized models on MATH-10K dataset. The fine-tuning hyper-parameters are show in Table 12.
Table 12: Hyper-parameters for the finetuning of Llama2-7B with quantized models.
Hyper-parameter Llama2-7B
number of blocks 64
Optimizer AdamW
Learning rate 7e-4
Weight decay 1.0
LR scheduler linear
Warmup ratio 10%
Epochs 3
Batch size 16
Max sequence length 512
Modules Q K V O G U D